OpenStudy (anonymous):

Find the quotient: 6x^3-5x^2+x+6/2x+3

OpenStudy (anonymous):

this is about how far i got 6x^3-5x^2+x+16/2x+3 2x+3/6x^3-5x^2+x+6 3x^2*2x=6x^2 6x^3+6x^2 6x^3-6x^3=0 -5x^2+6x^2=-11x^2 -11x^2+1=11x 11x(2x+3) -11x*2x+-11x*3 -22x+-33x

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

long division is very very hard to write here, but that is what you have to do

OpenStudy (anonymous):

if you need to see the steps worked out, i would suggest looking here http://www.wolframalpha.com/input/?i=%286x^3-5x^2%2Bx%2B6%29%2F%282x%2B3%29 go down to where it says "quotient and remainder" and click on "show steps" and you will see them worked out